i need help simplifying (x-2)(3x+5)
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

i need help simplifying
(x-2)(3x+5) = (x-2)(3x) + (x-2)5 = 3x^2-6x+5x-10 = 3x^2-x-10
=================
Cheers,
Stan H.
-----------------

Answer by rfer(16322)

3x^2+5x-6x-10=
3x^2-x-10